OpenStudy (anonymous):

Can anyone help me with this integral? im not sure what i am suppose to do with it

OpenStudy (anonymous):

\[\int\limits\limits_{0}^{infinity} e^(-\theta*x)(1/3)e^(-x/3)dx\]

OpenStudy (anonymous):

where theta >0

OpenStudy (hba):

Tried solving it ?

OpenStudy (anonymous):

ya i have the problem i come into is that i get different values for differnt thetas

OpenStudy (anonymous):

what makes sense i think so im not sure what im suppose to do with it

OpenStudy (hba):

\[\int\limits\limits\limits_{0}^{infty } e^{-\theta*x}(1/3)e^{-x/3}dx \] This is the correct ques?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

so i did a simmilar question which was the same thing but without the e^(-theta*x) part and i got 3 with i know is right

OpenStudy (anonymous):

int\limits\limits_{0}^{infinity} e^(-\theta*x)(1/3)e^(-x/3)dx theta >0 (apparantly)

OpenStudy (anonymous):

but there are only limits for x so how should i solve it?

OpenStudy (anonymous):

well it would be e^(-x^2/3) i think

OpenStudy (anonymous):

ya sorry that should be in there too

OpenStudy (anonymous):

so make x = ((-x*theta)/3)

OpenStudy (anonymous):

then im guessing after computing there will be a answer that is dependent on theta?

OpenStudy (anonymous):

well then its e^x with does not converge

OpenStudy (abb0t):

For infinite limits,you take the integral from [0, R] and then take the limit as R->∞

OpenStudy (abb0t):

sorry, for improper integrals* my mistake.

OpenStudy (abb0t):

Also note: \[\int\limits_{a}^{∞} \frac{ 1 }{ x^R }dx\] p>1 (convergent) p<1 (divergent)

OpenStudy (anonymous):

so how does that help me?

OpenStudy (abb0t):

idk. i'm just throwing random stuff out there I guess.

OpenStudy (hba):

lol